The forward and reverse mutation rates at a certain locus in E. coli are A to a 5 x 10-5 a to A 3 x 10-3 a. What are the expected allelic frequencies (p^ and q^) if no other processes are involved? b. If allele “a” is a deleterious recessive, will natural selection be able to easily eliminate it from this E. coli population? Why or why not?
Answer #1

Lets take Initial allele frequencies be:

  • p = freq (A) allele
  • q = freq (a) allele
  • Let u = rate per generation, for forward mutation when A changes to a = 5 x 10-5
  • Let v = rate per generation, for backward mutation when a mutates back to A = 3 x 10-3

At equilibrium, Frequency of A allele

p = v/u+v = 3*10^-3/5*10^-5 + 3*10^-3

= 0.9836

at equilibrium, frequency of allele

q = p = v/u+v = 5*10^-5/5*10^-5+ 3*10^-3

= 0.0164